Can someone please tell me why this is suddenly is an addition (indicated by the blue arrow) when the formula has a subtraction here.

Thanks in advance!!Exercise 25.2 The reaction 2 NO2(g) ? 2 NO(g) + O2(g) has a rate constant of k= 0.522 M-1 s-1 at 592 K, and a rate constant of k = 1.70 M-1 s-1 at 627 K. (a) Calculate Ea for the reaction, giving your answer in kJ mol-1. mal
Answer #1

Arhenius equation,

rate constant, k = Ae-Ea/RT

AT TWO DIFFERENT TEMPERATURES , DIFFERENT RATE CONSTANTS TO WRITE THE EQUATION TAKES AS,

In(K2/K1) = (Ea/R) [ (1/T1) -(1/T2] , THIS IS ONE ACTUAL CORRECT FORM AT TWO DIFFERENT TEMPERATURES.

EQUATION WRITTEN CORRECTLY. SUBSTITUTION ALSO CORRECT.

BUT WHAT IS WRONG?

SIGN WRONG.  

-1/627K IS TO BE WRITTEN

THE SOLUTION SHOWING +1/627K, WHICH IS WRONG.

2.303 log(1.70/0.522) = (Ea/8.314)[(1/592)-(1/627]

on solving above , Ea = (0.3655x 8.314 x 105) /9.4292

= 32227 joules

= 32.227 kj

--------------------------------

please note Inx means ,it is 2.303logx